OpenStudy (anonymous):

5.25 mL of an unknown HCl sample is titrated with 25.3 mL of a 0.00100 M solution of NaOH to its equivalence point, what is the initial concentration of the acid?

OpenStudy (anonymous):

@aaronq @sweetburger

OpenStudy (anonymous):

@seehearcreate

OpenStudy (anonymous):

Choices: 2.08*10^-2 M 4.82*10^-3 M 8.28*10^-4 M 4.64*10^-3 M

OpenStudy (anonymous):

I'd help you with this but I'm dealing with the same stuff so I'm not quite certain of myself haha!

OpenStudy (abmon98):

Construct a balanced equation NaOH+HCl-->NaCl+H2O Number of moles=Volume*concentration Number of moles of NaOH=(25.3/1000)*0.00100 Number of moles of NaOH=0.0000253 moles Ratio of moles of NaOH to HCl is 1:1 0.0000253=5.25/1000*concentration 1000*0.0000253/5.25=0.004819 when rounded we get 0.00482 so out of the choices i choose the second one

OpenStudy (anonymous):

Ok I was thinking the 2nd or 4th but I really had no idea so I'm gonna go with yours since you set it up. Thank you!!!

OpenStudy (anonymous):

And you're right :) Thank you SO much!

OpenStudy (abmon98):

Your most welcome :)